1) What should the period of a spring oscillator be if it has a mass of 5 kg and a spring constant of 25 N/m? What would the new period be if the mass is doubled?
Answer #1

Here\: in\; the\; question \; mass\; and \; spring \; constant\; is \; given:{\color{Blue} }

Formula \; of \; period \; of \; a \; spring \; oscillator\; in \; form \; of \; mass\; and \; spring \; constant\;

is\; below:

                                     T=2\pi \sqrt{M/K}

Now \; taking \; the \; first \; case \; when \; mass\; is =5kg

                                    T_{1}=2\pi \sqrt{5/25} \; \; \; \; \; ({\color{Blue} as\; the\; value\; of\; Spring\; constant\; is\; given \; K=25N/m)}

                                     T_{1}=2\pi \sqrt{1/5}\; second

                                      {\color{Red} T_{1}=2.808\; second}     (Ans)

{\color{Blue} Now \; take\; the \; seconed\; case\; when\; mass\; is\; doubled} :

                     In\; this\; case \; the \; mass \; will\; be\; 10kg.

                     So\; put\; the\; value \; of \; M\; and\; K \; in \; the\; below\; equation :                                                   T=2\pi \sqrt{M/K}

                                                      T_{2}=2\pi \sqrt{10/25}\; second

                                                       {\color{Red} T_{2}=3.97\; second}       (Ans)